OpenStudy (anonymous):

Can some help with question 2 and 4? I will post document.

OpenStudy (akshay_budhkar):

A] Propose a value for the following definite integral and give an interpretation for the result : is nothing but the area under the curve from the values x=6 to x=8

OpenStudy (saifoo.khan):

@Akshay_Budhkar , we can find that by simply using area of a square? right?

OpenStudy (akshay_budhkar):

yes 8^2=64 =D

OpenStudy (saifoo.khan):

Perfect. ;D

OpenStudy (saifoo.khan):

exactly..

OpenStudy (akshay_budhkar):

aha the y-axis is degrees / hour! not degrees!

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

It's not easy. haha

OpenStudy (akshay_budhkar):

u already did :P i am trying

OpenStudy (saifoo.khan):

haha.

OpenStudy (anonymous):

Amistre64 was helping me on this question but I was still lost so I was hoping if someone else also explained it maybe I would understand.

OpenStudy (akshay_budhkar):

ok i get it

OpenStudy (akshay_budhkar):

lets start with A

OpenStudy (akshay_budhkar):

are you clear with A ?

OpenStudy (saifoo.khan):

@Akshay_Budhkar idk but i feel like we have to find some derivatives?

OpenStudy (anonymous):

Yes I understand what you said with A.

OpenStudy (akshay_budhkar):

@saifoo.khan yes

OpenStudy (akshay_budhkar):

ok now let degrees be d and time be t so on y axis we have d/t and x axis we have t Now till x= 1 it is a straight line, with a slope -8 and intercept -8 so the equation is y=mx+c that is y=-8x-8 that is d/t=-8t-8 d=-8t^2=8t

OpenStudy (akshay_budhkar):

*d=-8t^2-8

OpenStudy (akshay_budhkar):

Now for minima or maxima the first derivative of the function is o therefore\[{d(d) \over dt }=0=-16t-8 \therefore t= 1/2\]

OpenStudy (akshay_budhkar):

Now to find if it is a minima or a maxima we take the second derivative. if it is negative it is a maxima if positive minima ( mostly i maybe mixing them ) so in this case it is -16 which is negative. therefore it is a maxima at 1/2

OpenStudy (anonymous):

Thank you so much I understood that! Would you be able to help me with question 4 by chance?

OpenStudy (akshay_budhkar):

ok lemme see, is it one unit for each square right?

OpenStudy (anonymous):

That is what I understand.

OpenStudy (akshay_budhkar):

ok lets start with the simplest C. f(3) = 5 any doubt?

OpenStudy (anonymous):

Yes, thats what I see.

OpenStudy (akshay_budhkar):

ok now A. f'(-2) Now i believe that at a given point \[f'(x)={dy \over dx} = { y \over x}\] if we estimate. so f'(-2) =-2/-2 = 1

OpenStudy (anonymous):

Okay, that makes sense to me.

OpenStudy (akshay_budhkar):

for D its the same logic, area under the curve. i believe you can assume the graph to be a triangle and calculate the area

OpenStudy (akshay_budhkar):

is it unclear?

OpenStudy (anonymous):

So you think D would be =-2 to 4?

OpenStudy (akshay_budhkar):

the area under the graph , from -2 to 4

OpenStudy (anonymous):

Okay

OpenStudy (akshay_budhkar):

Can you solve it? For the other questions i would personally plot a graph for f'(x)

OpenStudy (anonymous):

f"(x)=d/dx*f'(x)

OpenStudy (akshay_budhkar):

yes i know. Now to plot the graph u must know that it is zero at minima and maxima, and u can fidn approximate values by using the approximation of y/x, and plot the graph of f'

OpenStudy (akshay_budhkar):

then after i get the graph, f"' will be obtained by using y/x again and the last one by area under curve in that graph

OpenStudy (anonymous):

Sorry I am just thinking. Okay, Thanks.

OpenStudy (akshay_budhkar):

yup give it a thought, and lemme know if ustuck =D and u can tag me in some other question if u get some similar doubt =D

OpenStudy (anonymous):

Okay, thank you so much for your help!